OpenStudy (anonymous):

what is the solution set for 4x>8?

OpenStudy (anonymous):

4x>8 To get x alone, divide both sides by 4: \[\frac{ 4 }{ 4 }x=\frac{ 8 }{ 4 }\] Can you find x now? :o)

OpenStudy (anonymous):

Correction: \[\frac{ 4 }{ 4 }x>\frac{ 8 }{ 4 }\]